OpenStudy (anonymous):

how to simplify (x^3-8)/x-2

OpenStudy (akshay_budhkar):

i felt this was ishaan94!!

OpenStudy (anonymous):

\[x^3-8=(x-2)(x^2+2x+4)\] then cancel

OpenStudy (akshay_budhkar):

x^3-8=(x-2)(x^2+2x+4) so cancel x-2

OpenStudy (anonymous):

the top facotrs to (x-2)(x^2+2x+4) which means the (x-2) terms cancel out. Leaving only (x^2+2x+4)

OpenStudy (anonymous):

difference of two cubes \[a^3-b^3=(a-b)(a^2+ab+b^2)\]

OpenStudy (anonymous):

i mean the (x-2) terms divide out

OpenStudy (akshay_budhkar):

you get it?

OpenStudy (anonymous):

the answer is x^2-6

OpenStudy (anonymous):

no

OpenStudy (akshay_budhkar):

no you can see that answer is x^2+2x+4 as x-2 cancels out

OpenStudy (anonymous):

i will write it out \[\frac{x^3-8}{x-2}=\frac{(x-2)(x^2+2x+4)}{x-2}=x^2+2x+4\]

OpenStudy (anonymous):

@shana would you like to try it with numbers and see?